OpenStudy (anonymous):

Ingrid has a previous balance of $643 on a credit card with an 18.3% APR compounded monthly. If she made a payment of $50 this month, what is the new balance on her credit card?

OpenStudy (anonymous):

Her monthly rate would be 18.3/12 = 1.525% If she had 643 balance previously, this means that this months balance will be 643*1.01525= 652.81 (original amt plus interest) She makes a $50 payment, so remaining balance is 652.81-50 = 602.81.